OpenStudy (anonymous):

Please i need help!!!!! Medal and Fan! 1) 2^(3-y)= 8 2) 1/27=b^-3 3)1/81= k^-4

OpenStudy (anonymous):

hmm is it pretty clear that \(2^3=8\)?

OpenStudy (triciaal):

2)b^-3 = 1/b^3 1/27 = b^-3 = 1/b^3 1/27^(1/3) = 1/b^(3*1/3) 27^1/3 = b=

OpenStudy (triciaal):

3) 81 can be written as 3^4

OpenStudy (butterflydreamer):